A call option with a strike price of $78 on a stock selling at $97 costs $21.5. What are the call option’s intrinsic and time values? what are the Intrinsic value and Time value?
Answer #1

Intrinsic Value of Call Option = Max(0, (S - X))

So,

Intrinsic Value of Call Option = (97 - 78)

Intrinsic Value of Call Option = $19

Value of Call Option = Intrinsic Value + Time Value

So,

Time Value of Call Option = 21.50 - 19

Time Value of Call Option = $2.50